40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Nov 12 01:09:56 2016, gautam, Update, LSC, Recovering DRMI locking 
    Reply  Wed Nov 16 03:10:01 2016, gautam, Update, LSC, DRMI locked on 1f and 3f signals 
       Reply  Wed Nov 16 08:14:43 2016, Steve, Update, LSC, DRMI locked on 1f and 3f signals 5hrs.png
       Reply  Mon Nov 21 14:02:32 2016, gautam, Update, LSC, DRMI locked on 3f signals, arms held on ALS DRMIArms_Nov20.pdf
    Reply  Wed Nov 23 16:21:02 2016, gautam, Update, LSC, ITMY UL glitches are back ITMYproblemsBack.png
       Reply  Mon Nov 28 10:27:13 2016, gautam, Update, SUS, ITMY UL glitches are back ITMY_satboxSpectra.pdfITMY_UL_testerBox.png
       Reply  Tue Nov 29 11:07:37 2016, Steve, Update, LSC, ITMY UL glitches are back gliching400d.png
          Reply  Wed Nov 30 01:47:56 2016, gautam, Update, LSC, Suspension woes IMCwoes.png
             Reply  Wed Nov 30 17:08:56 2016, gautam, Update, LSC, Binary output breakout box removed 
                Reply  Thu Dec 1 02:19:13 2016, gautam, Update, LSC, Binary output breakout box restored MC1_glitch_fast.pngITMY_testerGlitch.pngITMYsatbox.pngULcomparison.pdf
                   Reply  Thu Dec 1 08:02:57 2016, Steve, Update, LSC, glitching ITMY_UL_LL ITMY_UL_LL.png
                      Reply  Fri Dec 2 11:56:42 2016, gautam, Update, LSC, MC1 LEMO jiggled IMG_3474.JPGMC1_LR_dead.png
                         Reply  Mon Dec 5 15:05:37 2016, gautam, Update, LSC, MC1 glitches are back Mc1glitches.png
                      Reply  Thu Dec 8 10:13:43 2016, Steve, Update, LSC, glitching ITMY_UL has a history glitching__ITMY-UL_2007.png
Message ID: 12657     Entry time: Fri Dec 2 11:56:42 2016     In reply to: 12654     Reply to this: 12664
Author: gautam 
Type: Update 
Category: LSC 
Subject: MC1 LEMO jiggled 

I noticed 2 periods of frequent IMC locklosses on the StripTool trace, and so checked the MC1 PD readout channels to see if there were any coincident glitches. Turns out there wasnt yes BUT - the LR and UR signals had changed significantly over the last couple of days, which is when I've been working at 1X5. The fast LR readback was actually showing ~0, but the slow monitor channel had been steady so I suspected some cabling shenanigans.

Turns out, the problem was that the LEMO connector on the front of the MC1 whitening board had gotten jiggled ever so slightly - I re-jiggled it till the LR fast channel registered similar number of counts to the other channels. All looks good for now. For good measure, I checked the 3 day trend for the fast PD readback for all 8 SOS optics (40 channels in all, I didn't look at the ETMs as their whitening boards are at the ends), and everything looks okay... This while situation seems very precarious to me, perhaps we should have a more robust signal routing from the OSEMs to the DAQ that is more immune to cable touching etc...

ELOG V3.1.3-